an electric motor rotating a workshop grinding wheel at 1.04x10^2 rev/min is switched off. assume the wheel gas a constant negative angular acceleration of magnitude 1.92 rad/s^2. a) how long does it take the grinding wheel to stop and b) through how many radians has the wheel turned during the tine interval found in part a)?
Answer #1

a)

t = w/a

t = (104* 2* 3.14 / 60) / 1.92

t = 5.67 s

======

b)

d = w^2/ ( 2 a)

d = (104* 2 * 3.14 / 60)^2 / ( 2* 1.92)

d = 30.857 rad

======

Comment before rate in case any doubt, will reply for sure.. goodluck